# Trivariate Bicycle Codes

### v1

@@ Abstract
Quantum error correction suppresses noise in quantum systems to allow for high-precision computations. In this work, we introduce Trivariate Bicycle Quantum Low-Density Parity-Check (TB-QLDPC) codes, via an extension of the framework developed by Bravyi et al. [Nature, 627, 778-782 (2024)]. Unlike the weight-6 codes proposed in their study, our approach also offers weight-4 and weight-5 codes, which promises to be more amenable to near-term experimental setups. We show that our TB-QLDPC codes up to weight-6 have a bi-planar structure. Further, most of our new codes can also be arranged in a two-dimensional toric layout, and have substantially better encoding rates than comparable surface codes while offering comparable error suppression capabilities. For example, we can encode $4$ logical qubits with distance $5$ into $30$ physical qubits with weight-5 check measurements, while a surface code with comparable parameters requires $100$ physical qubits. The high encoding rate and compact layout make our codes highly suitable candidates for near-term hardware implementations, paving the way for a realizable quantum error correction protocol.
I Introduction
@@
Quantum computing promises to solve problems intractable for classical computers [1, 2, 3]. Despite advancements in improving hardware quality, noise remains the most significant challenge for realizing practical implementations to ultimately achieve both reliable and scalable quantum computations. As a consequence, quantum error correction (QEC) is required to reduce noise of the physical hardware [4, 5]. Among many QEC codes [6, 7, 8], the surface code has emerged which currently leads the forefront in experimental quantum computing due to its relatively high threshold error rates [9, 8, 10, 11, 12, 13] and suitability for implementation with current hardware devices [14]. Despite its advantages, the surface code requires high qubit overhead leading to an asymptotically zero encoding rate while the distance scales as the square-root of the number of physical qubits [15].
Recently, Quantum Low-Density Parity-Check (QLDPC) codes have attracted attention for their potential to require less physical qubits than surface codes [16]. LDPC codes are well-established in classical error correction [17, 18, 19] for their efficiency and decoding performance. In a recent breakthrough [20], Bravyi et al. presented a set of QLDPC codes that are significantly more qubit-efficient than the surface code of the same number of logical qubits. These codes, called Bivariate Bicycle (BB) QLDPC codes, have origins in the work by Kovalev and Pryadko [21]. Ref. [20] introduced codes with weight-$6$ stabilizer checks which can be encoded into a bi-planar toric layout. This layout allows their code to accommodate four local stabilizer checks and two long-range checks. Still, finding codes that reduce the weight of the stabilizer checks while maintaining good code properties could significantly simplify hardware implementations of QLPDC codes.
@@
Here, we introduce trivariate bicycle (TB) QLDPC codes to provide several novel low-weight codes with better encoding rates than comparable surface codes. We define TB-QLDPC codes by introducing an additional variable in the generating polynomials of the BB-QLDPC framework. The additional variable increases the feasibility region of our optimization over code parameters and allows for more codes within the framework. As a highlight, we find a weight-$5$ code with a single long-range check which encodes $4$ logical qubits into only $30$ physical qubits in a bi-planar toric architecture, while a comparable surface code requires $100$ qubits. We analyze the performance of our codes under depolarizing noise, where we find comparable noise suppression to similar surface codes. We make progress on the properties of TB-QLPDC codes. We prove that our TB-QLDPC codes up to weight-$6$ possess a bi-planar structure (see Proposition 6 in Appendix C). This is not obvious given the presence of an extra variable and a different number of terms in the parity check matrices. Furthermore, we provide a weight-independent criterion to check whether a TB-QLDPC code has a toric structure (see Proposition 2). We also show that all stabilizer checks are translational-invariant when the TB-QLDPC codes has a toric layout (see Proposition 8). Finally, we show that weight-$4$ TB-QLDPC codes can possess a tangled toric layout, where the Tanner graph corresponds to a deformed torus with some mismatched edges.
II Background
QEC codes are essential for executing high-fidelity computations on noisy quantum computers. Without QEC, errors accumulate during circuit execution, yielding an unreliable output. Stabilizer codes [22, 23] represent a class of QEC codes defined by their stabilizer group, which is an abelian subgroup of the Pauli group on $n$ qubits that keeps the codespace invariant. The codespace of a stabilizer code is equivalently described as the joint $+1$-eigenspace of the stabilizer generators $S=\langle S_{1},S_{2},...,S_{r}\rangle$. For a $[[n,k,d]]$ code with $n$ physical qubits, $k$ logical qubits, and distance $d$, the number of linearly independent generators is $r=n\!-\!k$. A stabilizer code qualifies as a QLDPC code if each qubit supports a constant number of stabilizer generators and each generator has a constant weight independent of $n$. A stabilizer code is classified as a CSS code [24, 7] if each generator is a tensor product of $X$ and $I$ or a tensor product of $Z$ and $I$. Although the surface code is a QLDPC code, its encoding rate $k/n$ approaches zero as $n$ approaches infinity, contributing to its high overhead. In contrast, alternative QLDPC codes maintain asymptotically constant encoding rates while preserving or improving the $d=\mathcal{O}(\sqrt{n})$ distance scaling of the surface code [25, 26, 16, 27, 28, 29, 30].
III Framework
@@
Figure 1: Our weight-5 $[[30,4,5]]$ TB-LDPC code. It can be arranged in a bi-planar grid with periodic boundary conditions (PBC) using $30$ data qubits (which belong either to the L or R sublattice) and $30$ qubits for parity check measurements. Each parity check involves four neighbouring data qubits, and one non-local interaction described by a translational invariant lattice vector $(x,y)$. For X-checks, the interaction vector is $(-4,-3)$, while for Z-checks we have $(2,3)$. For better visibility, we only plot representatives of the non-local interactions at the center of the grid.
@@
Here, we introduce TB-QLDPC codes, which are a generalisation of BB-QLDPC codes discussed in Ref. [39, 20]. We start with the identity matrix $I_{m}$ and the cyclic shift matrix $S_{l}$ of size $l\times l$ with $(S_{l})_{ij}=\delta_{j,(i+1)\,\text{mod}\,l}$ for integer-valued $l$ and $m$. Let now
$x=S_{l}\otimes I_{m}\hskip 21.33955pty=I_{l}\otimes S_{m}\hskip 21.33955ptz=S_{l}\otimes S_{m}\,.$
be matrices of dimension $(l\cdot m)\!\times\!(l\cdot m)$ that commute with each other $[x,y]=0,[x,z]=0,[y,z]=0$. Previous BB-QLDPC codes are limited by two only variables, namely $x$ and $y$. Here, we introduce TB-QLDPC codes which are characterised by adding the third variable $z$ of the form $S_{l}\otimes S_{m}$. Now, we define the pair of matrices
@@
leading to a code of weight $\mathcal{W}=\mathcal{W}_{A}+\mathcal{W}_{B}$ with each matrix $A_{j}$ and ${B}_{j}$ being powers of $x,y$ or $z$. Working in the binary field $\mathbb{F}_{2}=\{0,1\}$, we ensure avoiding cancellation of terms by only allowing for unique terms in both $A$ and $B$. In order to fully describe a TB code, we need to define the stabilizer check matrices
@@
Following this definition, we can further define $X$-type and $Z$-type check operators being a row $v\!\in\!\mathbb{F}_{2}$ of $H_{X}$ and $H_{Z}$, respectively. Since $[A,B]=0$, these satisfy the stabilizer code condition requiring all $X$-type checks to commute with all $Z$-type checks, translating to $H_{X}H_{Z}^{T}=0$. We emphasize that the $A_{j}$ and ${B}_{j}$ terms can be understood as introducing edges between checks and vertices on the Tanner Graph of the TB-QLDPC code, one per term. This viewpoint is crucial in understanding physical implementations in Section V. Further details can be found in Appendix C.
We then find the TB-QLDPC code with parameters $[[n,k,d]]$ in accordance with Lemma 1 of Ref. [20]
@@
Figure 2: Comparison of noise properties between TB-QLDPC codes of weight-$4$ and weight-$5$. We find the fitted code distances $d_{\text{fit}}=\{5.9,12.0,6.1,8.3\}$, respectively.
Figure 3: Comparison of noise properties between small TB-QLDPC codes of of weight-$6$ and weight-$7$. We find the fitted code distances $d_{\text{fit}}=\{4.1,6.4,6.1\}$, respectively.
IV Results
@@
In Table 1 we provide an overview of our TB-QLDPC codes of weight five to seven and include some weight-4 TB-QDPC codes as well. For a given distance $d$ and number of logical qubits $k$, our codes require less physical qubits than surface codes which need $n=kd^{2}$ physical qubits [9, 8]. Note that to encode $k$ logical qubits via surface codes, one uses $k$ rotated planar surface codes with $d^{2}$ qubits each [20]. In Fig. 2, we illustrate decoding results under depolarizing noise for codes of stabilizer weights four and five and in Fig. 3 for weights six and seven. We performed decoding for physical error rates $p$ down to $10^{-3}$ and then extrapolate for lower error rates following a fit with Eq. (4).
@@
In Table 2 of Appendix A, we present an overview of additional QLDPC codes that outperform the surface code characteristics and provide a summary for the properties for codes of weight four to seven. We prove in Proposition 6 that our QLDPC codes of weight-$4$ and weight-$5$ allow for a bi-planar encoding architecture. This makes them suitable candidates for near-term hardware implementations.
$[[n,k,d]]$ $\mathcal{W}$ $r=k/n$ $r_{\text{BB}}/r_{\text{SC}}$ $p_{0}$ $p_{\text{L}}(10^{-4})$ toric bi-planar
$[[112,8,5]]$ 4 $1/14$ 1.8 0.0298 $2\!\times\!10^{-9}$$\times$ $\checkmark$
$[[144,2,12]]$ 4 $1/72$ 2.0 0.1017 $4\!\times\!10^{-19}$$\times$ $\checkmark$
$[[30,4,5]]$ 5 $1/7$ 3.3 0.0437 $6\!\times\!10^{-10}$checkmark $\checkmark$
$[[96,4,8]]$ 5 $1/24$ 2.7 0.0823 $1\!\times\!10^{-13}$checkmark $\checkmark$
$[[30,6,4]]$ 6 $1/5$ 3.2 0.0234 $3\!\times\!10^{-7}$checkmark $\checkmark$
$[[48,4,6]]$ 6 $1/12$ 3.0 0.0698 $3\!\times\!10^{-11}$checkmark $\checkmark$
$[[30,4,5]]$ 7 $1/7$ 3.3 0.0507 $5\!\times\!10^{-10}$checkmark unknown
@@
Table 1: Selection of TB-QLDPC codes grouped by weight $\mathcal{W}$ of the stabilizer checks. The codes parameters $[[n,k,d]]$ correspond to $k$ logical qubits, $n$ physical qubits and code distance $d$. The encoding rate $r=k/n$ is rounded to the next smallest integer value. In column $r_{\text{TB}}/r_{\text{SC}}$, we show the improvement factor in encoding rate compared to a surface code with same code distance and number of logical qubits. The last two columns provide the pseudo-threshold $p_{0}$ and the logical error $p_{\text{L}}$ at the physical error rate $p=10^{-4}$ which we gain from a fit with Eq. (4). We use a depolarizing error model without readout error.
@@
To benchmark our codes, we compare against planar surface code with similar distance $d$ and number of logical qubits $k$. In particular, the planar surface code of distance $d$ can encode one logical qubit into a $d\!\times\!d$ grid of physical qubits, leading to code parameters $[[d^{2},1,d]]$ [14]. To encode $k$ logical qubits, we implement $k$ instances of the planar surface code, resulting in a $[[kd^{2},k,d]]$ code. The logical error probability per syndrome cycle $p_{L}$ for $k$ logical qubits can be directly computed from the logical error of a single planar surface code $p_{L}(1)$ via [20]
@@
IV.1 Weight-4
In Fig. 4, we show the physical error rate $p$ against the logical error rate $p_{\text{L}}$ for our weight-4 QLDPC code $[[112,8,5]]$. As comparison, we also show surface codes with the same number of logical qubits and similar distance $d$. We find that our code has comparable error suppression than the $[[200,8,5]]$ surface code, while requiring only about half the number of physical qubits. We also fit the logical error rates with Eq. (4), where for our code with find a fitted code distance $d_{\text{fit}}=5.9$.
Figure 4: We show logical error rate $p_{\text{L}}$ against physical error rate $p$ for our weight-4 QLDPC code $[[112,8,5]]$, and surface codes with similar code parameters. The error curves are fitted with Eq. (4), where we report a fitted code distance of $d_{\text{fit}}=5.9$.
IV.2 Weight-5
Figure 5: Logical error rate $p_{\text{L}}$ against physical error rate $p$ for our weight-4 QLDPC code $[[112,8,5]]$. We show surface codes with comparable code parameters as reference. We find $d_{\text{fit}}=6.1$ for our QLDPC code.
Next, in Fig. 5 we study the error suppression of our $[[30,4,5]]$ code, which is a weight-$5$ TB-QLDPC codes. The layout of the code is shown in Fig. 1. Our code suppresses logical errors at a physical error rate $p=10^{-3}$ by more than three orders of magnitude. Our code allows for a toric bi-planar layout and has about three times higher encoding rate compared to the $[[100,4,5]]$ surface code which has similar error suppression.
IV.3 Weight-6
In Fig. 6, we present our weight-6 $[[30,6,4]]$ code which has high encoding rate of $1/5$. Further, we find that the code has a toric layout with bi-planar structure. We find comparable error suppression scaling as the distance $3$ surface code.
Figure 6: Logical error rate of our weight-6 QLDPC code $[[30,6,4]]$ against comparable surface codes. We find $d_{\text{fit}}=4.1$.
At physical error rate $p=10^{-3}$, we find a logical error rate of $3.5\times 10^{-5}$, which is a logical noise suppression by two orders of magnitude. We note that a surface code with distance-$4$ and same number of logical qubits requires $96$ physical qubits, making our code three times more qubit efficient.
IV.4 Weight 7
Now, we study our weight-7 code with parameters $[[30,4,5]]$ in Fig. 7. Note that we also found a weight-$5$ code with the the same code parameters. We find that our weight-7 $[[30,4,5]]$ code has similar error suppression than the $[[100,4,5]]$ surface code.
Figure 7: Our weight-7 QLDPC code $[[30,4,5]]$ compared against surface codes with same number of logical qubits. We find $d_{\text{fit}}=6.1$.
IV.5 Comparison with other QLDPC codes
In Fig. 8, we compare the noise suppression characteristics of our weight-$5$ $[[30,4,5]]$ TB-QLDPC code (Fig. 1) with the weight-6 $[[72,12,6]]$ code of Ref. [20] and the weight-6 $[[72,8,6]]$ code of Ref. [41]. As a code of distance $d$ can correct up to $\lfloor(d-1)/2\rfloor$ errors, all three codes can correct at most $2$ errors. All three codes can be encoded into a bi-planar two-dimensional grid with toric layout. We note that to encode $8$ logical qubits, one can take two independent copies of our $[[30,4,5]]$ code to get a weight-$5$ $[[60,8,5]]$ code. We highlight that this way, our code has a better encoding rate than the weight-$6$ code $[[72,8,6]]$ of Ref. [41]. Further, our code has only weight-$5$ parity checks and requires only one long-range connection, compared to the weight-$6$ checks with two long-range connections required by the codes of Refs. [20, 41].
Figure 8: Logical error rate of our weight-5 $[[30,4,5]]$ code, weight-6 $[[72,8,6]]$ code of Ref. [41], and weight-6 $[[72,12,6]]$ code of Ref. [20].
V Code layout
Here, we discuss the physical layout of qubits needed to implement our TB-QLDPC codes. We study two independent requirements, bi-planarity and a toric layout.
@@ V Code layout
All TB-QLDPC codes of weight-4, where $A=A_{1}+A_{3}$, B = $B_{1}+B_{3}$, all codes of weight-5 such that A = $A_{1}+A_{3}$, B = $B_{1}+B_{2}+B_{3}$, and all codes of weight-6 such that A = $A_{1}+A_{2}+A_{3}$, B = $B_{1}+B_{2}+B_{3}$ (the case presented in  [20]), or A = $A_{1}+A_{3}$, B = $B_{1}+B_{2}+B_{3}+B_{4}$ allow for a bi-planar architecture of thickness $\theta=2$. The bi-planar decomposition can be computed in time $O(n)$.
The proof is found in Appendix C Proposition 6. It is currently unknown whether the weight-7 code is bi-planar.
Beyond bi-planarity, we further characterize the Tanner Graph structure using the concept of a toric layout to simplify the design process in physical implementations as elaborated in Appendix C Proposition 5. We view the Tanner Graph as embedded on a torus, allowing for edge crossings.
@@
Not all our found codes satisfy this property of a toric layout. However, we should keep in mind that a toric layout is desirable, but not necessarily indispensable in hardware implementation.
Note that any Cayley Graph is connected since its edges are produced by its generators, and any toric layout code must have a subgraph isomorphic to a Cayley Graph. Thus, codes with multiple connected components, such as the weight-4 [[112, 8, 5]] code, cannot have a toric layout. However, we note that the code has four connected graph isomorphic components. This implies that there are four [[28, 2, 5]] CSS subcodes which are equivalent to one another, since each subgraph uniquely defines the Tanner Graph of a smaller CSS code.
The problem of whether a graph $G$ has a subgraph isomorphic to another graph $H$ is well known to be NP-complete [42]. However, using the following proposition for TB-QLDPC codes we obtain an easy check for the spanning subgraph problem. Let $\text{ord}(A)$ denote the order of an element $A$ in a group.
@@
Note that the above proposition is independent of the bi-planarity of the Tanner Graph in Proposition 6. The proof for the case $\mathcal{W}_{A}=\mathcal{W}_{B}=3$ is found in Ref. [20].
To explain the idea behind Proposition 2, split our data qubits into two blocks through the definition
$H_{X}=[A|B]\quad\text{and}\quad H_{Z}=[B^{T}|A^{T}]$
. L represents first half of the qubits (left side of block matrix), R represents the second half (right side of block matrix).
[…]
Although Ref. [20] proves Proposition 2 explicitly for the weight-6 case where $A=A_{1}+A_{2}+A_{3}$, $B=B_{1}+B_{2}+B_{3}$ for $i,j,g,h\in\{1,2,3\}$, we emphasize that the proposition holds more generally because none of their argument requires that $i,j,g,h\in\{1,2,3\}$. In fact, we may carry out an identical argument for generic weights $A=\sum^{\mathcal{W}_{A}}_{j=1}A_{j}$ and $B=\sum^{\mathcal{W}_{B}}_{j=1}B_{j}$. As an example of the generality of Proposition 2, our weight-7 $[[30,4,5]]$ code was numerically confirmed to have a toric layout for $\mu=l=5,\lambda=m=3$ and $i=1,j=g=2$, $h=4$.
We find that our weight-5 $[[30,4,5]]$ and $[[96,4,8]]$ codes have a toric layout, while the $[[72,4,8]]$ code does not. For example, the weight-5 $[[96,4,8]]$ code has $\mu=l=8,\lambda=m=6$ and i=g=1, j=h=2, h=4. Our weight-6 $[[48,4,6]]$ and weight-7 $[[30,4,5]]$ codes have a toric layout, with $2$ and $3$ long-range connections out of each vertex respectively.

### v2

# Multivariate Bicycle Codes

@@ Abstract
Quantum error correction suppresses noise in quantum systems to allow for high-precision computations. In this work, we introduce Multivariate Bicycle (MB) Quantum Low-Density Parity-Check (QLDPC) codes, via an extension of the framework developed by Bravyi et al. [Nature, 627, 778-782 (2024)] and particularly focus on Trivariate Bicycle (TB) codes. Unlike the weight-6 codes proposed in their study, we offer concrete examples of weight-4 and weight-5 TB-QLDPC codes which promise to be more amenable to near-term experimental setups. We show that our TB-QLDPC codes up to weight-6 have a bi-planar structure. Further, most of our new codes can also be arranged in a two-dimensional toric layout, and have substantially better encoding rates than comparable surface codes while offering similar error suppression capabilities. For example, we can encode $4$ logical qubits with distance $5$ into $30$ physical qubits with weight-5 check measurements, while a surface code with these parameters requires $100$ physical qubits. The high encoding rate and compact layout make our codes highly suitable candidates for near-term hardware implementations, paving the way for a realizable quantum error correction protocol.
@@
Introduction.—  Quantum computing promises to solve problems intractable for classical computers [1, 2, 3]. Despite advancements in improving hardware quality, noise remains the most significant challenge for realizing practical implementations to ultimately achieve both reliable and scalable quantum computations. As a consequence, quantum error correction (QEC) is required to reduce noise of the physical hardware [4, 5]. Among many QEC codes [6, 7, 8], the surface code has emerged which currently leads the forefront in experimental quantum computing due to its relatively high threshold error rates [9, 8, 10, 11, 12, 13] and suitability for implementation with current hardware devices [14]. Despite its advantages, the surface code requires high qubit overhead leading to an asymptotically zero encoding rate while the distance scales as the square-root of the number of physical qubits [15].
@@
Recently, Quantum Low-Density Parity-Check (QLDPC) codes have attracted attention for their potential to require less physical qubits than surface codes [16]. LDPC codes are well-established in classical error correction [17, 18, 19] for their efficiency and decoding performance. In a recent breakthrough [20], Bravyi et al. presented a set of QLDPC codes that are significantly more qubit-efficient than the surface code of the same number of logical qubits. These codes, called Bivariate Bicycle (BB) QLDPC codes, have origins in the work by Kovalev and Pryadko [21]. Ref. [20] introduced codes with weight-$6$ stabilizer checks which can be encoded into a bi-planar toric layout. This layout allows their code to accommodate four local stabilizer checks and two long-range checks. Still, finding codes that reduce the weight of the stabilizer checks while maintaining good code properties could significantly simplify hardware implementations of QLPDC codes. Here, we introduce Multivariate Bicycle (MB) QLDPC codes to provide several novel low-weight codes with better encoding rates than comparable surface codes. We define MB-QLDPC codes in a general manner, and then focus on the concrete case of Trivariate Bicycle (TB) codes, which have three variable in the generating polynomials. For improved readability, we will from now on only use the term code to refer to a QLDPC code, if not explicitly stated otherwise. As a highlight, we find a weight-$5$ code with a single long-range check which encodes $4$ logical qubits into only $30$ physical qubits in a bi-planar toric architecture, while a comparable surface code would require $100$ qubits. We analyze the performance of our codes under depolarizing noise, where we find comparable noise suppression to similar surface codes. We make progress on the properties of TB codes and prove that our codes up to weight-$6$ possess a bi-planar structure in Proposition 6 of SM F. This is not obvious given the choice of terms in the parity check matrices. Furthermore, we provide a weight-independent criterion to check whether a TB code has a toric layout structure in Proposition 2. We also show that all stabilizer checks are translational-invariant when the MB-QLDPC codes has a toric layout, further details to be found in Proposition 9. Finally, we show that weight-$4$ TB codes can possess a tangled toric layout, where the Tanner graph corresponds to a deformed torus with some mismatched edges.
Framework.—  The group algebra of the multivariate bicycle code with $r$ variables is $G_{r}=\nicefrac{{\mathbb{Z}}}{{l_{1}}}\times\nicefrac{{\mathbb{Z}}}{{l_{2}}}\times\cdots\times\nicefrac{{\mathbb{Z}}}{{l_{r}}}$. In the special case of $r=2$, we recover the group algebra of the BB code.
@@
Figure 1: Our weight-5 $\llbracket 30,4,5\rrbracket$ TB-QLDPC code can be arranged in a bi-planar grid with periodic boundary conditions (PBC) using $30$ data qubits (which belong either to the L or R sublattice) and $30$ qubits for parity check measurements. Each parity check involves four neighbouring data qubits, and one non-local interaction described by a translational invariant lattice vector $(x,y)$. For X-checks, the interaction vector is $(-4,-3)$, while for Z-checks we have $(2,3)$. For better visibility, we only plot representatives of the non-local interactions at the center of the grid.
The group $G_{r}$ is isomorphic to the quotient of the multivariate polynomial ring $R_{r}=\nicefrac{{\mathbb{F}_{2}\left[x_{1},x_{2},\cdots,x_{r}\right]}}{{\left(x_{1}^{l_{1}}-1,x_{2}^{l_{2}}-1,\cdots,x_{r}^{l_{r}}-1\right).}}$ Here, the quotient ring $R_{r}$ has a basis consisting of the monomials $x_{1}^{i_{1}}x_{2}^{i_{2}}\cdots x_{r}^{i_{r}}$, where $10\leq i_{1}\leq l_{1}-1,0\leq i_{2}\leq l_{2}-1,\cdots,0\leq i_{r}\leq l_{r}-$ 1 . In this work, we define the variable $x_{i}$ using the following encoding:
  1. 1.
    Let $enc(i)=b_{q}b_{q-1}\cdots b_{2}b_{1}$ be the binary encoding of $i$ using $q=\lfloor\log r\rfloor+1$ bits.
  2. 2.
    Let $t_{\alpha}\equiv S_{l_{i}}^{b_{\alpha}}$ where $S_{l_{i}}$ is a cyclic shift matrix of size $l_{i}\times l_{i}$.
  3. 3.
    $x_{i}\equiv t_{q}\otimes t_{q-1}\otimes\cdots t_{2}\otimes t_{1}.$
Note that the above encoding is not unique and has been chosen for its simplicity and one can trivially define other encodings. Moreover, the different variables are not independent. For example, in the trivariate case, the quotient ring is given by
$R_{3}=\nicefrac{{\mathbb{F}_{2}\left[x_{1},x_{2},x_{3}\right]}}{{\left(x_{1}^{l_{1}}-1,x_{2}^{l_{2}}-1,x_{3}-x_{1}x_{2}\right).}}$
@@
In general, one can define variables such that they are all independent. For concreteness, in this work, we will focus on $m=3$ which will make them TB-QLDPC codes, a generalisation of BB-QLDPC codes discussed in Ref. [22, 20]. We start with the identity matrix $I_{m}$ and the cyclic shift matrix $S_{l}$ of size $l\times l$ with $(S_{l})_{ij}=\delta_{j,(i+1)\,\text{mod}\,l}$ for integer-valued $l$ and $m$. Previous BB codes are limited by two only variables, namely $x$ and $y$. Here, we introduce TB codes which are characterised by adding the third variable $z$ of the form $S_{l}\otimes S_{m}$. Let now
$x\equiv x_{1}=S_{l}\otimes I_{m}\,,\hskip 5.69046pty\equiv x_{2}=I_{l}\otimes S_{m}\,,\hskip 5.69046ptz\equiv x_{3}=S_{l}\otimes S_{m}$
be matrices of dimension $(l\cdot m)\!\times\!(l\cdot m)$ that commute with each other $[x_{i},x_{j}]=0$ $\forall\,i,j\!\in\!\{1,2,3\}$. Now, we define the pair of matrices
@@
generating a code of weight $\mathcal{W}=\mathcal{W}_{A}+\mathcal{W}_{B}$ with each matrix $A_{j}$ and ${B}_{j}$ being powers of $x,y$ or $z$. Working in the binary field $\mathbb{F}_{2}=\{0,1\}$, we ensure avoiding cancellation of terms by only allowing for unique terms in both $A$ and $B$. In order to fully describe a TB code, we need to further define the stabilizer check matrices
@@
Following this definition, we can then define $X$-type and $Z$-type check operators being a row $v\!\in\!\mathbb{F}_{2}$ of $H_{X}$ and $H_{Z}$, respectively. Since $[A,B]=0$, these satisfy the stabilizer code condition requiring all $X$-type checks to commute with all $Z$-type checks, translating to $H_{X}H_{Z}^{T}=0$. We emphasize that the $A_{j}$ and ${B}_{j}$ terms can be understood as introducing edges between checks and vertices on the Tanner Graph of the TB code, one per term. This viewpoint is crucial in understanding physical implementations in the Code Layout section. Further details can be found in SM F.
We then find the TB code with parameters $\llbracket n,k,d\rrbracket$ in accordance with Lemma 1 of Ref. [20]
@@
Figure 2: Comparison of noise properties among our TB codes of weight-$4$ and weight-$5$. We find the fitted code distances $d_{\text{fit}}=\{5.9,12.0,6.1,8.3\}$, respectively.
@@
Results.—  In Table 1, we provide an overview of our TB codes of weights four to seven. For a given distance $d$ and number of logical qubits $k$, our codes require less physical qubits than surface codes which need $n=kd^{2}$ physical qubits [9, 8]. Note that to encode $k$ logical qubits via surface codes, one uses $k$ rotated planar surface codes with $d^{2}$ qubits each [20]. In Fig. 2, we illustrate decoding results under depolarizing noise for codes of stabilizer weights four and five. We performed decoding for physical error rates $p$ down to $10^{-3}$ and then extrapolate for lower error rates following a fit with Eq. (4).
@@
In Table 2 in the SM, we present an overview of additional TB codes that outperform the surface code characteristics and provide a summary for the properties for codes of weights four to seven. We prove in Proposition 6 that our codes of weight-$4$ and weight-$5$ allow for a bi-planar encoding architecture. This makes them suitable candidates for near-term hardware implementations.
$\llbracket n,k,d\rrbracket$ $\mathcal{W}$ $r=k/n$ $r_{\text{TB}}/r_{\text{SC}}$ $p_{0}$ $p_{\text{L}}(10^{-4})$ toric bi-planar
$\llbracket 144,2,12\rrbracket$ 4 $1/72$ 2.0 0.1017 $4\!\times\!10^{-19}$$\times$ $\checkmark$
$\llbracket 30,4,5\rrbracket$ 5 $1/7$ 3.3 0.0437 $6\!\times\!10^{-10}$checkmark $\checkmark$
$\llbracket 30,6,4\rrbracket$ 6 $1/5$ 3.2 0.0234 $3\!\times\!10^{-7}$checkmark $\checkmark$
$\llbracket 30,4,5\rrbracket$ 7 $1/7$ 3.3 0.0507 $5\!\times\!10^{-10}$checkmark unknown
@@
Table 1: Selection of TB-QLDPC codes ordered by weight $\mathcal{W}$ of the stabilizer checks. The code parameters $\llbracket n,k,d\rrbracket$ correspond to $k$ logical qubits, $n$ physical qubits and code distance $d$. The encoding rate $r=k/n$ is rounded to the next smallest integer value. In column $r_{\text{TB}}/r_{\text{SC}}$, we show the improvement factor in encoding rate compared to a surface code with the same code distance and number of logical qubits. The last two columns provide the pseudo-threshold $p_{0}$ and the logical error $p_{\text{L}}$ at the physical error rate at $p=10^{-4}$ which we gain from a fit with Eq. (4). We use a depolarizing error model without readout error.
@@
For benchmarking, we compare our codes against planar surface codes with similar distance $d$ and number of logical qubits $k$. In particular, the planar surface code of distance $d$ can encode one logical qubit into a $d\!\times\!d$ grid of physical qubits, leading to code parameters $\llbracket d^{2},1,d\rrbracket$ [14]. To encode $k$ logical qubits, we implement $k$ instances of the planar surface code, resulting in a $\llbracket kd^{2},k,d\rrbracket$ code. The logical error probability per syndrome cycle $p_{L}$ for $k$ logical qubits can be directly computed from the logical error of a single planar surface code $p_{L}(1)$ via [20]
@@
In Fig. 3, we compare the noise suppression characteristics of our weight-$5$ $\llbracket 30,4,5\rrbracket$ TB code (Fig. 1) with the weight-6 $\llbracket 72,12,6\rrbracket$ code of Ref. [20] and the weight-6 $\llbracket 72,8,6\rrbracket$ code of Ref. [24]. As a code of distance $d$ can correct up to $\lfloor(d-1)/2\rfloor$ errors, all three codes can correct at most $2$ errors. Also, all of them encoded into a bi-planar two-dimensional grid with toric layout. We note that to encode $8$ logical qubits, one can take two independent copies of our $\llbracket 30,4,5\rrbracket$ TB code to get a weight-$5$ $\llbracket 60,8,5\rrbracket$ code. We highlight that this way, our TB code has a better encoding rate than the weight-$6$ BB code $\llbracket 72,8,6\rrbracket$ of Ref. [24]. Further, our code has only weight-$5$ parity checks and requires only one long-range connection, compared to the weight-$6$ checks with two long-range connections required by the codes of Refs. [20, 24].
Figure 3: Logical error rate of our weight-5 $\llbracket 30,4,5\rrbracket$ TB code, weight-6 $\llbracket 72,8,6\rrbracket$ BB code of Ref. [24], and weight-6 $\llbracket 72,12,6\rrbracket$ BB code of Ref. [20].
Code Layout.—  Here, we discuss the physical layout of qubits needed to implement our TB-QLDPC codes. We study two independent requirements, bi-planarity and a toric layout.
@@
All TB-QLDPC codes of weight-4, where $A=A_{1}+A_{3}$, B = $B_{1}+B_{3}$, all codes of weight-5 such that A = $A_{1}+A_{3}$, B = $B_{1}+B_{2}+B_{3}$, and all codes of weight-6 such that A = $A_{1}+A_{2}+A_{3}$, B = $B_{1}+B_{2}+B_{3}$ (the case presented in Ref. [20]), or A = $A_{1}+A_{3}$, B = $B_{1}+B_{2}+B_{3}+B_{4}$ allow for a bi-planar architecture of thickness $\theta=2$. The bi-planar decomposition can be computed in time $O(n)$.
The proof is found in Proposition 6. It is currently unknown whether the weight-7 code is bi-planar.
Beyond bi-planarity, we use the concept of a toric layout to simplify the design process in physical implementations as elaborated in Proposition 5. We view the Tanner Graph as embedded on a torus, allowing for edge crossings, see Fig. 1.
@@
Using the proposition below, we obtain an easy check for whether a TB code has a toric layout.
@@
We discuss further intuition in Ref. 8.
Although Ref. [20] proves Proposition 2 for the weight-6 case where $A=A_{1}+A_{2}+A_{3}$, $B=B_{1}+B_{2}+B_{3}$, we emphasize that the proposition holds more generally because none of their argument requires that $i,j,g,h\in\{1,2,3\}$. In fact, we may carry out an identical argument for generic weights $A=\sum^{\mathcal{W}_{A}}_{j=1}A_{j}$ and $B=\sum^{\mathcal{W}_{B}}_{j=1}B_{j}$. As an example of the generality of Proposition 2, our weight-7 $\llbracket 30,4,5\rrbracket$ code was numerically confirmed to have a toric layout, see Table 2.
Note that since $r=2$ for our code, our toric layout criterion and results are on 2-dimensional Euclidean space which is experimentally relevant. If our multivariate group algebra was comprised was $r$ independent variables, we obtain an analogous toric layout definition and criterion on an $r$-dimensional torus for $\mathbb{Z}_{2\mu_{1}}\times...\times\mathbb{Z}_{2\mu_{r}}$, which experiment is not ready to realize. This higher dimensional toric layout arises because each independent variable introduces an independent orthogonal direction in Euclidean space.